OpenStudy (anonymous):

(Geometry) If D-M-G, and if DM = 2x, MG = x-5, and DG = & then find x. Use Betweeness of Points or Rays.

OpenStudy (anonymous):

DG=7

OpenStudy (anonymous):

DG=DM+MG 7=2x+x-5 2=3x 2/3=x

OpenStudy (anonymous):

I thought so too, but when I checked the answer it was x = 4.

OpenStudy (anonymous):

ohh i see, i made a mistake 7=2x+x-5 7+5=2x+x 12=3x 4=x

OpenStudy (anonymous):

Ahhh thank you so much.